# Reducing Hallucinations in Language Model-based SPARQL Query Generation Using Post-Generation Memory Retrieval

### v2

@@ Abstract
The ability to generate SPARQL queries from natural language questions is crucial for ensuring efficient and accurate retrieval of structured data from knowledge graphs (KG). While large language models (LLMs) have been widely adopted for SPARQL query generation, they are often susceptible to hallucinations and out-of-distribution errors when producing KG elements like Uniform Resource Identifiers (URIs) based on internal parametric knowledge. This often results in content that appears plausible but is factually incorrect, posing significant challenges for their use in real-world information retrieval (IR) applications. This has led to increased research aimed at detecting and mitigating such errors. In this paper, we introduce PGMR (Post-Generation Memory Retrieval), a modular framework that incorporates a non-parametric memory module to retrieve KG elements and enhance LLM-based SPARQL query generation. Our experimental results indicate that PGMR consistently delivers strong performance across diverse datasets, data distributions, and LLMs. Notably, PGMR significantly mitigates URI hallucinations, nearly eliminating the problem in several scenarios.
@@
Aditya Sharma1,2, Luis Lara2, Christopher J. Pal1,2,3, Amal Zouaq1,2 1Polytechnique Montréal, 2Mila - Quebec AI Institute, 3Canada CIFAR AI Chair Correspondence: aditya.sharma@mila.quebec
@@ 1 Introduction
Question answering (QA) is a critical task in natural language processing and remains a key goal for artificial intelligence research. Many QA methodologies employ knowledge graphs (KGs) (Saxena et al., 2020; Lan et al., 2021; Peng et al., 2023), which are multi-relational graphs that encode facts. KGs represent world knowledge as fact triples (subject, relation, object), for example, (The Truman Show, nominated for, Academy Award for Best Supporting Actor). The problem of answering questions that require reasoning over KGs (KGQA) can be broken down into two parts, namely, translating the natural language question into its associated SPARQL [^1] https://www.w3.org/TR/rdf-sparql-query/ query and then executing the SPARQL query over a KG to obtain the final answer (Banerjee et al., 2022). For example, a natural language question like "For which film was Sergei Eisenstein the film editor?" can be translated into the following SPARQL query - "select distinct ?sbj where { ?sbj wdt:p1040 wd:q8003 . ?sbj wdt:p31 wd:q11424 }," where p1040, q8003, p31 and q11424 are Uniform Resource Identifiers (URIs) (described further in Section 3.2) representing entities, relations and properties from a KG like Wikidata [^2] https://www.wikidata.org. Our approach specifically focuses on the following challenge: translating a question into the appropriate semantically accurate SPARQL query and ensuring the precise generation of URIs linked to the entities and relationships mentioned in the question.
Using large language models (LLMs) for generating SPARQL queries has become prevalent in recent research (Meyer et al., 2024; Diallo et al., 2024; Banerjee et al., 2022; Qi et al., 2024; Brei et al., 2024). However, this approach can lead to significant challenges, such as hallucinations and out-of-distribution errors, particularly when the models generate KG elements like URIs based only on their internal parametric knowledge. As a consequence of these hallucinations, SPARQL queries may be generated with a structure that seems correct while incorporating incorrect factual information, such as URIs that don’t exist in the KG (Huang et al., 2023; Ji et al., 2023; Huang et al., 2021). This limitation raises concerns about the reliability of LLMs in real-world information retrieval (IR) applications and has driven extensive research into techniques for detecting and mitigating such errors (Lin et al., 2024; Varshney et al., 2023; Dhuliawala et al., 2023; Chern et al., 2023; Li et al., 2023).
@@
Retrieval-augmented generation (RAG) (Khandelwal et al., 2019; Lewis et al., 2020a; Guu et al., 2020; Lewis et al., 2020b) is a popular paradigm that is often employed for reducing hallucinations (Huang et al., 2023). It uses a non-parametric memory for retrieving information in response to a question. This information is provided in the prompt for the LLM before generation to enhance the grounding of responses. Despite the integration of retrieval mechanisms, RAG often retrieves noisy or incomplete information. Additionally, since the LLM is not strictly constrained to rely on the retrieved context, it may still generate responses that are not fully grounded, thereby failing to eliminate hallucinations (Barnett et al., 2024).
In response to these limitations, we propose PGMR (Post-Generation Memory Retrieval), a new modular architecture featuring a non-parametric memory retriever module for managing KG elements. This approach enables LLMs to focus on generating SPARQL query syntax while the retriever ensures accurate retrieval of relevant KG elements. PGMR enables the LLM to generate intermediate SPARQL queries with natural language entities and relations enclosed in special tokens, which allows for controlled hallucination. These hallucinations are later corrected and grounded in KG URIs using the retriever. By isolating the tasks of query generation and element retrieval, PGMR improves the reliability and accuracy of SPARQL queries, avoiding the hallucination of non-existent KG URIs and reducing out-of-distribution errors, as discussed in Section 6.
In this context, we address the following research questions: (1) Does the integration of a memory retrieval module into an LLM-based system enhance performance in SPARQL query generation and mitigate hallucination issues? (2) Does the addition of a memory module increase model robustness to out-of-distribution evaluation?
@@
    We explore the implications of adding a post-generation memory retrieval module to LLMs for SPARQL query generation, investigating reduction in LLM hallucination and improvements in query accuracy and resilience against out-of-distribution scenarios.
@@
    We propose PGMR, an innovative modular architecture where initial SPARQL query syntax generation is performed by an LLM, followed by retrieval of KG elements from a non-parametric memory module.
@@
    We introduce a paradigm for letting LLMs hallucinate and fixing these hallucinations post-generation using a non-parametric memory.
@@
Figure 1: As discussed in Section 4.1, PGMR first employs an LLM to produce an intermediate query, where the URI labels are framed by "starturi" and "enduri" tokens (Section 4.1.1). The retriever (Section 4.1.2) subsequently fetches and replaces these labels with the most similar URIs from the non-parametric memory to create the final SPARQL query.
@@ 2 Related Work
SPARQL Query Generation using LLMs. Even though various models have been proposed for the SPARQL query generation task recently, they suffer from one key problem: they require tagged question data (Banerjee et al., 2022; Reyd and Zouaq, 2023; Diallo et al., 2024; Qi et al., 2024). In tagged data, the input to the model contains the natural language question along with the URIs and the URI labels needed to formulate the corresponding SPARQL query. This requirement presents a challenge due to the high costs and labor-intensive nature of tagging, making it unsuitable for practical, real-world use cases. Therefore, our study emphasizes and assesses the untagged versions of the datasets.
Reyd and Zouaq (2023) and Diallo et al. (2024) propose a copy mechanism for copying URIs from the tagged question while generating the SPARQL query using LLMs. Additionally, they investigate the models’ performance in dealing with novel question-query structures and unknown URIs. We incorporate this evaluation approach to evaluate the out-of-distribution robustness of our proposed approach. Banerjee et al. (2022) propose a pointer generator network-based approach for SPARQL query generation over tagged data using BERT (Kenton and Toutanova, 2019) and using a finetuned T5. Their approach generates multiple SPARQL queries and selects the first query that executes and fetches an answer from the knowledge base as the predicted SPARQL query. This evaluation strategy contrasts with our approach of generating only one SPARQL query per question, which is more efficient. Qi et al. (2024) employ a pre-training method that integrates the proposed Triple Structure Correction (TSC) strategy. This technique, which requires tagged data, involves randomly interchanging the positions of the subject, predicate, and object in SPARQL query triples with a certain probability alongside a Masked Language Modeling (MLM) objective in a multi-task learning setup before proceeding with finetuning on the downstream tagged dataset.
@@
External memory augmented LLMs. Earlier studies have investigated methods to augment LLMs by retrieving documents from external memory and integrating them into the contextual framework of tasks to provide relevant information. According to Modarressi et al. (2023) and Modarressi et al. (2024), the lack of a specialized memory unit in current LLMs constrains their ability to explicitly store and retrieve knowledge relevant to tasks. This observation supports our study’s central premise. They suggest enhancing LLMs with an API-based memory unit for read-write operations to address this limitation. Interactions with the external memory can be implemented using natural language interfaces (Park et al., 2023; Zhou et al., 2023) or formal mechanisms such as standardized APIs that allow the model to parse and execute commands (Schick et al., 2024; Hu et al., 2023; Modarressi et al., 2023, 2024). Our approach is akin to the use of a formal mechanism for memory retrieval through our intermediate query format (as described in Section 4.1.1).
@@ 3 Preliminaries
3.1 Knowledge Graphs (KGs)
A Knowledge Graph (KG) is a multi-relational graph with entities (e.g., "The Truman Show" and "Academy Award for Best Supporting Actor") as nodes and relations $r$ (e.g., "nominated for") as typed edges between the entity nodes {$s,o$}. Thus, each edge of the KG represents a fact {$s,r,o$}, e.g., {The Truman Show, nominated for, Academy Award for Best Supporting Actor}.
@@ 3.2 Uniform Resource Identifiers (URIs)
Each entity, relation, and property in a large KG like Wikidata is associated with a Uniform Resource Identifier or URI and a natural language URI label. For example, for an entity with the URI label "The Truman Show," the associated Wikidata URI is "q214801." There are two types of URIs in the Wikidata KG namely, "Q-ids" which correspond to the entities (e.g., "q76" refers to the entity "Barack Obama"), and "P-ids" which correspond to relations, properties, and classes in the KG (e.g., "p26" refers to the relation "spouse"). Generating Wikidata URIs within a SPARQL query using LLMs presents a significant challenge as each URI is just one letter followed by a random number. The LLM can only depend on its internal parametric memory to remember these URI mappings.
4 Proposed Model
In this section, we present our proposed model for SPARQL query generation using LLMs, termed PGMR (described in Section 4.1). Our model employs an external knowledge base, such as a KG, as a non-parametric memory to enhance the grounding of LLM generations. In PGMR, the LLM first generates an intermediate query with SPARQL syntax and KG elements in natural language (as discussed in Section 4.1.1) given only the question and then incorporates the KG URIs from memory post-generation to construct the SPARQL query.
4.1 Post-Generation Memory Retrieval (PGMR)
LLM-driven SPARQL query generation has emerged as a focus of recent research (Meyer et al., 2024; Diallo et al., 2024; Banerjee et al., 2022; Qi et al., 2024; Brei et al., 2024). Despite its potential, this approach is susceptible to significant issues, including hallucinations and out-of-distribution errors, particularly when KG elements like URIs are generated without direct grounding in external knowledge sources. Due to these hallucinations, LLM-generated SPARQL queries may maintain syntactic correctness while introducing factual errors, such as URIs that do not exist within the KG.
To address these issues, we introduce PGMR (“Post-Generation Memory Retrieval”), a novel method for SPARQL query generation utilizing LLMs. In PGMR, we initially convert SPARQL queries into an intermediate query format by substituting each URI with special "starturi" and "enduri" tokens, placing the natural language URI label between them, as outlined in Section 4.1.1. This modified data is then used to either finetune or few-shot the LLM. During inference, our retriever (discussed in Section 4.1.2) converts the intermediate queries generated by the LLM back into SPARQL queries by retrieving the URIs whose labels are the closest match in latent space to the predicted labels within the "starturi" and "enduri" markers in the intermediate query, as discussed in Section 4.1.3. PGMR’s approach of generating only one URI label between each pair of "starturi" and "enduri" tokens simplifies the retriever’s task, thereby increasing the accuracy of retrieving the correct URI. Unlike LLM-based direct SPARQL query generation, which may produce URIs not present in the KG, PGMR’s approach of conducting memory retrieval post-LLM generation mitigates the risk of hallucinating non-existent KG URIs, as discussed in Section 6.
Figure 2: To teach the LLM, the training data is first converted from (a) SPARQL to (b) intermediate queries, as detailed in Section 4.1.1
PGMR enables the LLM to generate intermediate SPARQL queries with natural language entities and relations enclosed in special tokens. These intermediate queries are later grounded in KG URIs using the retriever to produce the final SPARQL query. In PGMR, the LLM generates intermediate SPARQL queries containing natural language entities and relations wrapped in special tokens, which are later grounded in KG URIs through the use of a retriever to produce the final SPARQL query. To facilitate this, the training data is first transformed from the (a) SPARQL query to the (b) intermediate query to teach the LLM how to generate intermediate queries
Figure 3: The retriever finds each natural language URI label in the generated intermediate query using the special tokens and retrieves the corresponding URI from memory using similarity search, as discussed in Section 4.1.2
4.1.1 Data Transformation
Predicting the appropriate Wikidata URIs for SPARQL queries poses a significant challenge for LLMs because they lack pre-training on this specific data type. The models must rely solely on their parameters to recall mappings between the natural language descriptions of entities and relations in the question, and their corresponding URIs. In contrast, LLMs find it significantly easier to predict natural language labels for these URIs given a question. Based on this understanding, we transform the SPARQL queries in our training data into intermediate queries, thus simplifying the SPARQL query generation task. As seen in Figure 2, we substitute each URI in the original query with its label, which we obtain from Wikidata using its official Python library [^3] https://pypi.org/project/Wikidata/. We then surround these labels with "starturi" and "enduri" tokens, which our memory retriever (described in Section 4.1.2) uses to identify the URI labels’ locations.
4.1.2 Memory Retriever
Memory: We utilize a memory module designed as a dictionary that pairs LLM encoder-generated embeddings of URI labels and their descriptions as keys with the corresponding URIs as values. This module includes label embeddings and URIs for all URIs in the dataset.
Retriever: Given a natural language input such as "The Truman Show," our memory retriever locates the URI with the label most similar to this text within the memory. As illustrated in Figure 3, the text is initially encoded into an embedding by an encoder such as BGE (Chen et al., 2024) or MPNet (Song et al., 2020). The retriever then uses FAISS (Douze et al., 2024; Johnson et al., 2019) to retrieve the URI whose label embedding is nearest to the input text embedding from the memory. We use BGE as the encoder for our retriever in this study.
4.1.3 Setup
Training Setup: Using the transformed data from Section 4.1.1, we finetune the LLM to generate intermediate queries given the questions.
Few-shot Setup: Drawing on the transformed data, we apply few-shot prompting to the LLM, enabling it to generate intermediate queries corresponding to the input question.
Inference: During the inference phase, we initially produce an intermediate query for each question using the LLM. The intermediate query includes predicted URI labels marked by "starturi" and "enduri" tokens. Each of these labels is subsequently replaced by its corresponding URI through the retriever, thereby forming the final SPARQL query.
Model Type Query EM $\uparrow$ BLEU $\uparrow$ URI EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small finetuned 35.08 % 81.87 38.44 % 27.75 %

### v3

@@ Abstract
The ability to generate SPARQL queries from natural language questions is crucial for ensuring efficient and accurate retrieval of structured data from knowledge graphs (KG). While large language models (LLMs) have been widely adopted for SPARQL query generation, they are often susceptible to hallucinations and out-of-distribution errors when generating KG elements, such as Uniform Resource Identifiers (URIs), based on opaque internal parametric knowledge. We propose PGMR (Post-Generation Memory Retrieval), a modular framework where the LLM produces an intermediate query using natural language placeholders for URIs, and a non-parametric memory module is subsequently employed to retrieve and resolve the correct KG URIs. PGMR significantly enhances query correctness (SQM) across various LLMs, datasets, and distribution shifts, while achieving the near-complete suppression of URI hallucinations. Critically, we demonstrate PGMR’s superior safety and robustness: a retrieval confidence threshold enables PGMR to effectively refuse to answer queries that lack support, and the retriever proves highly resilient to memory noise, maintaining strong performance even when the non-parametric memory size is scaled up to 9 times with irrelevant, distracting entities.
@@
Aditya Sharma1,2, Christopher J. Pal1,2,3, Amal Zouaq1,2 1Polytechnique Montréal, 2Mila - Quebec AI Institute, 3Canada CIFAR AI Chair Correspondence: aditya.sharma@mila.quebec
@@ 1 Introduction
Question answering (QA) is a fundamental task in natural language processing research. Many QA systems leverage knowledge graphs (KGs) (Saxena et al., 2020; Lan et al., 2021; Peng et al., 2023), which represent world knowledge as multi-relational graphs of factual triples (subject, relation, object), for example, (The Truman Show, nominated for, Academy Award for Best Supporting Actor). Answering questions over KGs (KGQA) can typically involve two steps: first, translating a natural language question into its corresponding SPARQL [^1] https://www.w3.org/TR/rdf-sparql-query/ query (discussed in Section 3.2), and second, executing that query against the KG to retrieve an answer (Banerjee et al., 2022).
Large language models (LLMs) have shown promise for this task (Meyer et al., 2024; Diallo et al., 2024; Banerjee et al., 2022; Qi et al., 2024; Brei et al., 2024). However, the process of authoring SPARQL queries remains demanding for non-experts and is non-trivial even for modern LLMs, as it mandates an intimate familiarity with the underlying KG schema and the obscure structure of its Uniform Resource Identifiers (URIs - described in Section 3.1). Specifically, the use of non-semantic identifiers within platforms such as Wikidata (e.g., a prefix followed by arbitrary digits like Q937 for Albert Einstein) forces LLMs to rely solely on their parametric memory for accurate identifier mapping, impeding robust query generation. Furthermore, LLMs often hallucinate and produce plausible-looking but incorrect identifiers, such as URIs that don’t exist in the KG (Huang et al., 2023; Ji et al., 2023; Huang et al., 2021). This limitation raises concerns about the reliability of LLMs in real-world information retrieval (IR) applications and has driven extensive research into techniques for detecting and mitigating such errors (Lin et al., 2024; Varshney et al., 2023; Dhuliawala et al., 2023; Chern et al., 2023; Li et al., 2023).
@@
Retrieval-augmented generation (RAG) (Khandelwal et al., 2019; Lewis et al., 2020b; Guu et al., 2020; Lewis et al., 2020a) is a popular paradigm that is often employed for reducing hallucinations (Huang et al., 2023). It uses a non-parametric memory for retrieving information in response to a question. This information is provided in the prompt for the LLM before generation to enhance the grounding of responses. However, since the LLM is not strictly constrained to rely on the retrieved context, it may still generate responses that are not fully grounded, thereby failing to eliminate hallucinations (Barnett et al., 2024).
In response to these limitations, we propose PGMR (Post-Generation Memory Retrieval)  [^2] Code and data available at https://github.com/Lama-West/PGMR, a new modular architecture featuring non-parametric memory modules with a retriever for managing KG elements. This approach enables LLMs to focus on generating SPARQL query syntax while the retriever ensures accurate retrieval of relevant KG elements. PGMR enables the LLM to generate intermediate queries with a structured query representation that preserves SPARQL syntax while replacing KG URIs with natural language placeholders. Each placeholder is paired with the corresponding URI label and description in a mapping block. This design enables LLMs to focus on generating syntactically correct query structures, while deferring the grounding of placeholders to precise URIs to a post-generation retrieval step.
The core conceptual innovation of PGMR lies in its temporal repositioning of the retrieval phase. While RAG relies on pre-generation retrieval from natural language questions that often lack semantic alignment with URI metadata in the memory, PGMR performs retrieval after generation and leverages the LLM to generate explicit labels and descriptions after the query structure is formed. By producing these placeholders in a format that directly mirrors the non-parametric memory, PGMR achieves significantly higher latent similarity compared to RAG. This architecture effectively decouples structural synthesis from identifier grounding, ensuring that all generated URIs are strictly aligned with existing knowledge graph elements to nearly eliminate hallucinations.
We address the following research questions:
  1. 1.
    Does separating query structural prediction from identifier grounding reduce URI hallucinations while improving query correctness?
  2. 2.
    How effectively does a retrieval confidence threshold enable accurate query refusal under incomplete knowledge conditions?
  3. 3.
    How robust is PGMR’s performance when subjected to substantial memory scale-up and irrelevant data injection?
@@
    We propose PGMR, a modular architecture that enhances LLM-based SPARQL generation by explicitly separating query syntax (generated by the LLM) from URI resolution (handled by a non-parametric retriever).
@@
    We demonstrate that a retrieval confidence threshold enables PGMR to correctly refuse to answer queries based on incomplete knowledge, thereby ensuring safety and robustness.
@@
    PGMR exhibits minimal performance degradation (an approximate 4% SQM drop) when the memory is subjected to a ninefold increase in memory size and irrelevant data.
@@ 2 Related Work
SPARQL Query Generation using LLMs. Even though various models have been proposed for the SPARQL query generation task recently, they suffer from one key problem: they require tagged question data (Banerjee et al., 2022; Reyd and Zouaq, 2023; Diallo et al., 2024; Qi et al., 2024). In tagged data, the input to the model contains the natural language question, along with the URIs and their labels, necessary to formulate the corresponding SPARQL query. This requirement presents a challenge due to the high costs and labor-intensive nature of tagging, making it unsuitable for practical, real-world use cases. Therefore, our study emphasizes and assesses the untagged versions of the datasets.
Reyd and Zouaq (2023) and Diallo et al. (2024) propose a copy mechanism for copying URIs from the tagged question while generating the SPARQL query using LLMs. Additionally, they investigate the models’ performance in dealing with novel question-query structures and unknown URIs. We incorporate this evaluation method to measure the out-of-distribution robustness of our proposed approach. Banerjee et al. (2022) propose a pointer generator network-based approach for SPARQL query generation over tagged data, utilizing BERT (Kenton and Toutanova, 2019) and a finetuned T5. Their approach generates multiple SPARQL queries and selects the first query that executes and fetches an answer from the knowledge base as the predicted SPARQL query. This evaluation strategy contrasts with our approach of generating only one SPARQL query per question, which is more efficient. Qi et al. (2024) employ a pre-training method that integrates the proposed Triple Structure Correction (TSC) strategy. This technique, which requires tagged data, involves randomly interchanging the positions of the subject, predicate, and object in SPARQL query triples with a certain probability, alongside a Masked Language Modeling (MLM) objective in a multi-task learning setup, before proceeding with fine-tuning on the downstream tagged dataset. Zahera et al. (2024) explore chain-of-thought prompting (Wei et al., 2022) for SPARQL query generation. RAG for SPARQL query generation has also been explored in recent works (Emonet et al., 2024). Our comparative baseline adopts a comparable RAG approach.
@@
External memory augmented LLMs. Earlier studies have investigated methods to augment LLMs by retrieving documents from external memory and integrating them into the contextual framework of tasks to provide relevant information. According to Modarressi et al. (2023) and Modarressi et al. (2024), the lack of a specialized memory unit in current LLMs constrains their ability to store and retrieve knowledge relevant to tasks explicitly. This observation supports the central premise of our study. They suggest enhancing LLMs with an API-based memory unit for read-write operations to address this limitation. Interactions with external memory can be implemented using natural language interfaces (Park et al., 2023; Zhou et al., 2023) or formal mechanisms, such as standardized APIs, that allow the model to parse and execute commands (Schick et al., 2024; Hu et al., 2023; Modarressi et al., 2023, 2024). Our approach is akin to the use of a formal mechanism for memory retrieval through our intermediate query format (as described in Section 4.1).
@@ 3.1 Uniform Resource Identifiers (URIs)
In a large KG like Wikidata, each entity, relation, and property is assigned a Uniform Resource Identifier (URI), which is associated with a corresponding natural language URI label and other relevant metadata. For example, for an entity with the URI label "The Truman Show," the associated Wikidata URI is "Q214801." There are two types of URIs in the Wikidata KG, namely, "Q-ids," which correspond to the entities (e.g., "Q76" refers to the entity "Barack Obama"), and "P-ids," which correspond to relations, properties, and classes in the KG (e.g., "P26" refers to the relation "spouse").
3.2 SPARQL Queries
SPARQL is the standard declarative query language for Resource Description Framework (RDF) based KGs, such as Wikidata. It enables data retrieval by matching triple patterns corresponding to subject-predicate-object relationships. For example, a natural language question like "For which film was Sergei Eisenstein the film editor?" can be translated into the following SPARQL query - "select distinct ?sbj where { ?sbj wdt:p1040 wd:q8003 . ?sbj wdt:p31 wd:q11424 }," where p1040, q8003, p31, and q11424 are URIs representing entities, relations, and properties from a KG like Wikidata [^3] https://www.wikidata.org. However, authoring SPARQL queries is challenging for both non-experts and LLMs, as it requires precise knowledge of the KG’s schema and opaque URIs. Generating Wikidata’s non-semantic URIs (e.g., a letter followed by arbitrary digits) forces an LLM to rely solely on its parametric memory to recall identifier mappings.
Figure 1: As discussed in Section 4, PGMR first employs an LLM to produce an intermediate SPARQL-PGMR query, where all entities and relations are replaced by placeholders which are mapped to natural language labels and definitions, rather than formal URIs (Section 4.1). The retriever (Section 4.2) subsequently fetches and replaces these labels with the most similar URIs from the relevant non-parametric memory to create the final SPARQL query.
4 Post-Generation Memory Retrieval (PGMR)
In this section, we introduce PGMR (“Post-Generation Memory Retrieval”), a novel method for SPARQL query generation utilizing LLMs. In PGMR, we initially convert SPARQL queries into an intermediate SPARQL-PGMR query format. This intermediate query representation preserves the syntax of SPARQL while replacing URIs with placeholders that explicitly encode entities and relations in natural language. Specifically, entities are denoted as entityX = [ENT] Label [/ENT] Description, where Label provides the canonical surface form of the entity and Description gives a concise disambiguating context. Relations follow a parallel scheme, expressed as relationY = [REL] Label [/REL] Description, where the description clarifies the semantics of the property and distinguishes it from closely related relations. These intermediate queries are later grounded in KG URIs using a retriever (discussed in Section 4.2) to produce the final SPARQL query.
As seen in Figure 1, a SPARQL-PGMR intermediate query thus consists of two components: (1) a SPARQL query block that maintains the original query structure but replaces URIs with entity and relation placeholders, and (2) a mapping block that defines the corresponding natural language labels and disambiguating descriptions for each placeholder. This design ensures syntactic fidelity to SPARQL while maintaining human readability, thereby enabling LLMs to generate valid queries that can be reliably grounded to knowledge graph URIs in a post-processing step. This representation achieves two goals: (i) it constrains generation to a valid SPARQL structure, ensuring compatibility with downstream execution, and (ii) it provides natural language grounding for entities and relations, enabling post-generation alignment with the KG using a non-parametric memory of URIs and a retriever. By decoupling the SPARQL structural generation from the disambiguation of identifiers, SPARQL-PGMR supports accurate query generation and facilitates transparent inspection of model outputs. Unlike LLM-based direct SPARQL query generation, which may produce URIs not present in the KG, PGMR’s approach of conducting memory retrieval post-LLM generation greatly mitigates the risk of hallucinating non-existent KG URIs, as discussed in Section 6.
Model Type F1 $\uparrow$ SQM $\uparrow$ BLEU $\uparrow$ Qid EM $\uparrow$ Pid EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small FT 44.84 37.82 % 81.54 42.93 % 89.69 % 30.02 %
T5-Small RAG FT 67.30 67.03 % 89.96 73.68 % 95.48 % 14.70 %